OpenStudy (anonymous):

Simplify: 2√50 - 3√8

Parth (parthkohli):

\(\Large \color{Orange}{\Rightarrow 2\sqrt{50} = 2\sqrt{2 * 5 * 5} = 10\sqrt{2} }\) ====AND==== \(\Large \color{Orange}{\Rightarrow 3\sqrt{8} = 3\sqrt{2 * 2 * 2} = 6\sqrt{2} }\) Use the same property I told ya

Parth (parthkohli):

I'm getting 4 radical 2

OpenStudy (anonymous):

Can you write the property for me again

OpenStudy (anonymous):

How'd you get 4 radical 2?

Parth (parthkohli):

Okay \(\Large \color{Orange}{\Rightarrow x\sqrt{a} \pm y\sqrt{a} = (x \pm y)\sqrt{a} }\)